OpenStudy (anonymous):

x^4-3x^2+2=0 simplify, use integer or fraction and radicals as needed. express complex numbers in terms of i.

OpenStudy (dumbcow):

treat this like a quadratic let u = x^2, then u^2 = x^4 -> u^2 -3u +2 = 0 Factor -2*-1 = 2 -2+ -1 = -3 -> (u-2)(u-1) = 0 u = 2 u = 1 Now replace u with x^2 in order to solve for x x^2 = 2 --> x = +-sqrt(2) x^2 = 1 --> x = +-1

OpenStudy (anonymous):

I don't quite understand the answer's format. is that a plus and a minus in front of the sqrt of 2 and in front of the 1?

OpenStudy (dumbcow):

yes, since its a4th degree polynomial, there will be 4 roots \[x = 1,-1,\sqrt{2}, -\sqrt{2}\]

OpenStudy (anonymous):

ohhhh.... I get it now. thank you!